A particle is moving with a uniform acceleration along a straight line. Its velocities at A and B are respectively 7m/s^2 and 17m/s^2. M is the mid point of A and B. If t1is the time tajen to go from A to M and t2 is the time taken to go from M to B. Then what would be the ratio of t1/t2?

Grade:11

5 Answers

Adikesh Kumar
7 years ago
the qustion is not proper. Is it a acceleration or velocity? . if velocity then it must be in m/s 
repost correctly !!
 
aditya
7 years ago
dear student or sir 
a small mistake occured in your typing. velocity is considered in m/s and acceleration is considered in m/s^2. so I solve velocity in this question
given conditions in the question are
  1. particle is having uniform velocity.
  2. velocity at A is 7m/s
  3. velocity at B is 17m/s
  4. M is midpoint of AB. so let us consider the distance between AB is “d”
  5. therefore distance between AM is d/2 and distance between MB is d/2
  6. here t1 is the time taken to reach A to M. 
  7. so velocity=displscement /time
  8. time= displacement / velocity
  9. for case AM
  10. t1 =d/2*7  (*means multiplication)
  11. similarily t2=d/2*17
  12. t1/t2= 17/7 (on solving)
therefore the solution is t1/t2= 17/7 
SubWoof 330 X
3 years ago
Aditya your answer is too wrong its ans should b 3:2. I know the answer but I am yet to fing its solution😐😥
anonymous777 pathania
2 years ago
first we need to find velocity at midpoint M.let it be vms=(17^2 - 7^2 ) / 2as/2=( vm^2 - 49 ) / 2afrom these two equations we get vm= 13m/stime is directly proportional to (v-u) since a is constant ...implies that t1 is directly proportional to ( 13-7) i.e 6also t2 is directly proportional to (17-13) i.e 4from these two equations we get t1/ 12 = 6/4 i.e 3/2